# Model-based Deep Learning for Joint RIS Phase Shift Compression and WMMSE Beamforming

### v2

@@ III-B WMMSE Beamforming Updater
To solve (P1), we update $\textbf{W}_{\text{opt}}$ with $\boldsymbol{\theta}$ from the AQE, i.e. $\textbf{W}=f_{\text{w}}(\mathcal{I};\boldsymbol{\theta})$. To that end, we use the WMMSE algorithm [16] on the effective channel with the RIS phase shifts:
@@ G $=\textbf{H}_{AU}+\textbf{H}_{RU}\operatorname{diag}(e^{j\boldsymbol{\theta}})\textbf{H}_{AR}$ (5)
where $\textbf{G}^{T}=[\textbf{g}_{1},\ldots,\textbf{g}_{K}]^{T}\in\mathbb{C}^{M\times K}$ for all $K$ UEs given $\boldsymbol{\theta}$. The WMMSE solution for the AP beamforming vector of the $k$-th UE from [16, eq. (15)] is:
@@
where $\textbf{u}=[u_{1},\ldots,u_{K}]\in\mathbb{C}^{K}$, and $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{K}]\in\mathbb{R}_{+}^{K},\lambda_{k}\geq 0$ are the WMMSE receiver and weight parameters, obtained iteratively using an alternating optimization approach given $\textbf{W}_{\text{opt}}$ as an initial value.
To train the AQE-WMMSE end-to-end, we unroll the iterative algorithm in [16] into the WMMSE beamforming updater depicted in Figure 2. We first propose the NN: $\textbf{w}^{\prime}=f_{w^{\prime}}(\textbf{W}_{\text{opt}},\textbf{G})$ as an initialization stage, where the vector $\textbf{w}^{\prime}\in\mathbb{R}^{2MK}$ represents updating the optimal beamforming matrix $\textbf{W}_{\text{opt}}$ based on the actual channel G. Then we propose the NN: $\textbf{u},\boldsymbol{\lambda}=f_{u\lambda}(\textbf{w}^{\prime},\textbf{G})$ to learn the WMMSE receiver and weight parameters [16, eq. (5) and (13)]. All together, $f_{w^{\prime}}$ initializes the beamformer, while $f_{u\lambda}$ and (6) implement a single iteration of the unrolled WMMSE algorithm (i.e. repeating $f_{u\lambda}$ and (6) as additional layers in the DL architecture could replicate multiple iterations of [16, Table 1]). Finally, the WMMSE beamforming matrix is $\textbf{W}=\frac{\sqrt{P}}{\sqrt{\sum_{k=1}^{K}\|\textbf{w}_{k}\|^{2}}}[\textbf{w}_{1},\ldots,\textbf{w}_{K}]\in\mathbb{C}^{M\times K}$ which satisfies the transmit power constraint.
@@ IV-A DL Parameters
$f_{\text{EN}}$: Similar to [5], denoting $H=N+2MK$, we use a fully connected (FC) NN consisting of five linear layers of $32H$, $16H$, $8H$, $4H$, and $N_{c}$ neurons, respectively. The first four layers use the rectified linear unit (ReLU) activation function, while the last layer has no activation function. Between each pair of consecutive linear layers, we apply batch normalization followed by dropout with a probability of $0.5$.
$f_{\text{Q}}$: The scalar quantizer is set to 1-bit, i.e., $D=2$. Consequently, each encoder output $\psi_{n}$ corresponds to one bit of the control message, making the total number of control bits $B=N_{c}$. The amplitude $a_{1}$ is a non-trainable parameter set to $\frac{\pi}{2}$ resulting in two quantization levels: $\pm\frac{\pi}{2}$. The trainable parameters are initialized as $b_{1}=0$, and $c_{1}=0.5$.
$f_{\text{DE}}$: The decoder [^5] Since the decoder NN is implemented at the RIS controller, the decoder NN is designed to require less computation than the encoder NN at the AP. We also note that it is more difficult to update the parameters in the decoder than in the encoder, due to the vanishing gradients when passing information through the flat regions of the quantization layer [13]. is an FC NN with three linear layers each having $N$ neurons. The first two layers use ReLU, while the output layer has no activation function.
$f_{w^{\prime}}$ and $f_{u\lambda}$: both are four-layer FC NNs. Each of the first three layers has $4MK$ neurons (equal to the number of inputs) with a ReLU activation function followed by a dropout layer ($0.5$ probability). The last layer of $f_{w^{\prime}}$ and $f_{u\lambda}$ comprise $2MK$ and $3K$ neurons, respectively. No activation function is applied to all $2MK$ outputs of $f_{w^{\prime}}$ and $2K$ outputs of $f_{u\lambda}$, while the absolute value function is applied to the remaining $K$ outputs of $f_{u\lambda}$ to ensure non-negative elements for $\boldsymbol{\lambda}$.
@@ IV-B Training Procedure
We generate a dataset of size $100,000$ and use $64,000$, $16,000$, and $20,000$ samples for training, validation, and testing, respectively. In all experiments, we use a batch size of $L=128$, and an early stopping strategy with patience of 50 to reduce overfitting, with the maximum number of epochs set to $1,000$. The learning rate is initialized to $0.001$ and is reduced by a factor of $0.8$ whenever the validation loss does not decrease for 20 consecutive epochs, until it reaches a minimum value of $0.00005$. The training loss function for all models is:
${\mathcal{L}}(\textbf{W},\boldsymbol{\theta})=-\frac{1}{S}\sum_{l=1}^{L}\sum_{k=1}^{K}p_{k}\mathcal{R}_{k}$ (7)
where $S$ is the sample size. For fair comparison to previous work, we use a priority $p_{k}=1$ for all $K$ UEs.
@@ IV-C Benchmarks for Comparison
Upper Bound: we use the WMMSE-PI algorithm [4] with 100 iterations, to obtain the optimal AP beamforming matrix $\textbf{W}_{\text{opt}}$ and continuous valued RIS phase shifts $\boldsymbol{\theta}_{\text{opt}}$.
Benchmark 1 (DQNN [5]): a DL network that quantizes the phase shifts to $\pm\frac{\pi}{2}$, with each quantized value encoded using one bit. Therefore, the resulting control message has length $B=N$. The same DL network is also used to obtain the beamforming matrix based on the CSI [5, eq. (10), Fig. 2].
Benchmark 2 (ACFNet [6]): a DL network that compresses the bits of the control message. The DNN layer from [5, Fig. 2] is used as the input of the ACFNet method with $32H$, $16H$, $8H$, $4H$, and $2M+N$ neurons ($H=N+2MK$), in [6, Fig. 2] to obtain the beamforming matrix and phase shift vector before compression. For a fair comparison, compression is set to $B=N_{c}$ control bits by removing the Policy Network in [6, eq. (8) and Fig. 2] and the phase shifts at the RIS are continuous valued by removing $Q_{2}$ in [6, eq. (9) and Fig. 2].
Benchmark 3 (linQ [14]): a linear quantization method that updates the beamformer and compresses the bits of the control message. Given $\textbf{W}_{\text{opt}}$ and $\boldsymbol{\theta}_{\text{opt}}$, we implement linear analog combining and quantization [14] (i.e., NNs are matrix multiplication with bias and no activation function) to solve (P1), where $\boldsymbol{\theta}=f_{\text{D}}(f_{\text{Q}}(f_{\text{C}}(\boldsymbol{\theta}_{\text{opt}})))$ and $\textbf{W}=f_{W}(\textbf{W}_{\text{opt}})$. The linear operations ($f_{\text{C}}$, $f_{\text{D}}$, $f_{W}$) compress/decompress the phase shifts and update the beamformer.
All benchmarks use the same quantization scheme $f_{\text{Q}}(\cdot)$ outlined in Section IV-A to construct the control message. In the DQNN, ACFNet, and linQ benchmark methods, the beamforming matrix W is not calculated as a function of the actual phase shifts $\boldsymbol{\theta}$ chosen by the RIS controller.
@@ IV-D Results
In Fig. 3, we compare the achievable sum-rate for the proposed method with and without the WMMSE beamforming updater module (AQE-WMMSE and AQE) against the benchmarks. In Figure 3(a), we show the performance when the number of bits is fixed to be the number of RIS elements $B=N_{c}=N=100$. We also present the performance of each method when the control message fails to reach the RIS controller, in which case the RIS uses random phase shifts. Figures 3(b) and 3(c) compare the compression performance of each method where the number of bits of the control message becomes lower than the number of RIS elements, $B=N_{c}\leq N=100$. We can see that for high transmit power, AQE-WMMSE maintains a high achievable sum rate when compressing down from 100 bits to 40 bits and outperforms other methods even when compressing down to 10 bits.
@@ V Conclusion
A wireless communication informed DL architecture was proposed to jointly compress the RIS phase shifts and update the AP beamforming matrix by reconstructing the phase shift information into a control message (transmitted between the AP and RIS controller). Numerical results showed that updating the beamforming matrix based on the effective channel (with the actual reflection coefficients of the RIS) obtained a higher achievable sum rate than recent literature methods when compressing the number of bits in the control message to be lower than the number of RIS elements.

### v3

@@ III-B WMMSE Beamforming Updater
To solve (P1), we update $\textbf{W}_{\text{opt}}$ with $\boldsymbol{\theta}$ from the AQE, i.e. $\textbf{W}=f_{\text{w}}(\mathcal{I};\boldsymbol{\theta})$. To that end, we use the WMMSE algorithm [Shi2011] on the effective channel with the RIS phase shifts:
@@ G $=\textbf{H}_{AU}+\textbf{H}_{RU}\operatorname{diag}(e^{j\boldsymbol{\theta}})\textbf{H}_{AR}$ (5)
where $\textbf{G}^{T}=[\textbf{g}_{1},\ldots,\textbf{g}_{K}]^{T}\in\mathbb{C}^{M\times K}$ for all $K$ UEs given $\boldsymbol{\theta}$. The WMMSE solution for the AP beamforming vector of the $k$-th UE, adapted from [Shi2011, eq. (15)] to the case of a single-antenna UE, by dropping the Lagrange multiplier term $\mu_{k}$ and replacing it with the reciprocal SNR term $\sigma^{2}/P$, is:
@@
where $\textbf{u}=[u_{1},\ldots,u_{K}]\in\mathbb{C}^{K}$, and $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{K}]\in\mathbb{R}_{+}^{K},\lambda_{k}\geq 0$ are the WMMSE receiver and weight parameters, obtained through alternating optimizations, given $\textbf{W}_{\text{opt}}$ as the initial value. The WMMSE receivers help mitigate UE interference where the estimated data symbol at the $k$-th UE is then $\hat{s}_{k}=u_{k}^{*}y_{k}$ where $(\cdot)^{*}$ denotes complex conjugate.
To train the AQE-WMMSE end-to-end, we unroll the iterative algorithm in [Shi2011] into the WMMSE beamforming updater depicted in Fig. 2. We first propose the NN: $\textbf{w}^{\prime}=f_{w^{\prime}}(\textbf{W}_{\text{opt}},\textbf{G})$ as an initialization stage, where the vector $\textbf{w}^{\prime}\in\mathbb{R}^{2MK}$ represents updating the optimal beamforming matrix $\textbf{W}_{\text{opt}}$ based on the actual channel G. Then we propose the NN: $\textbf{u},\boldsymbol{\lambda}=f_{u\lambda}(\textbf{w}^{\prime},\textbf{G})$ to learn the WMMSE receiver and weight parameters [Shi2011, eq. (5) and (13)]. All together, $f_{w^{\prime}}$ initializes the beamformer, while $f_{u\lambda}$ and (6) implement a single iteration of the unrolled WMMSE algorithm. Finally, the WMMSE beamforming matrix is $\textbf{W}=\frac{\sqrt{P}}{\sqrt{\sum_{k=1}^{K}\|\textbf{w}_{k}\|^{2}}}[\textbf{w}_{1},\ldots,\textbf{w}_{K}]\in\mathbb{C}^{M\times K}$ which satisfies the transmit power constraint.
@@ IV-A DL Parameters
$f_{\text{EN}}$: Similar to [Xu2022], denoting $H=N+2MK$, we use a fully connected (FC) NN consisting of five linear layers of $32H$, $16H$, $8H$, $4H$, and $N_{c}$ neurons, respectively. The first four layers use the rectified linear unit (ReLU) activation function, while the last layer has no activation function. Between each pair of consecutive linear layers, we apply batch normalization followed by dropout with a probability of $0.5$.
$f_{\text{Q}}$: The scalar quantizer is set to 1-bit, i.e., $D=2$. Consequently, each encoder output $\psi_{n}$ corresponds to one bit of the control message, making the total number of control bits $B=N_{c}$. The amplitude $a_{1}$ is set to $\frac{\pi}{2}$ as a non-trainable parameter, resulting in two quantization levels: $\pm\frac{\pi}{2}$. The trainable parameters are initialized as $b_{1}=0$ and $c_{1}=0.5$.
$f_{\text{DE}}$: The decoder [^5] Since the decoder NN is implemented at the RIS controller, it is designed to be less complex than the encoder NN at the AP. We also note that it is easier to update the parameters in the decoder than in the encoder, due to the vanishing gradients from the flat regions of the quantization layer [Goodfellow2016a, Shohat2019]. is an FC NN with three linear layers each having $N$ neurons. The first two layers use ReLU, while the output layer has no activation function.
$f_{w^{\prime}}$ and $f_{u\lambda}$: both are four-layer FC NNs. Each of the first three layers has $4MK$ neurons (equal to the number of inputs) with a ReLU activation function followed by a dropout layer ($0.5$ probability). The last layer of $f_{w^{\prime}}$ and $f_{u\lambda}$ comprise $2MK$ neurons for $\textbf{w}^{\prime}$ and $3K$ neurons for $(\textbf{u},\boldsymbol{\lambda})$, respectively. No activation function is applied to all outputs of $f_{w^{\prime}}$. The first $2K$ outputs of $f_{u\lambda}$ for u use no activation function, while the absolute value function is applied to the remaining $K$ outputs of $f_{u\lambda}$ to ensure non-negative real valued elements for $\boldsymbol{\lambda}$.
@@ IV-B Training Procedure
We generate a dataset of size $100,000$ Monte Carlo simulations and use $64,000$, $16,000$, and $20,000$ samples for training, validation, and testing, respectively. In all experiments, we use a batch size of $L=128$, an early stopping strategy with patience of 50 to reduce overfitting, with the maximum number of epochs set to $1,000$. The learning rate is initialized to $0.001$ and reduced by a factor of $0.8$ when the validation loss does not decrease for 20 consecutive epochs, until reaching a minimum value of $0.00005$. The loss function being minimized for each batch is: ${\mathcal{L}}(\textbf{W},\boldsymbol{\theta})=-\sum_{l=1}^{L}\sum_{k=1}^{K}p_{k}\mathcal{R}_{k}$.
For fair comparison to previous work, we use a priority $p_{k}=1$ for all $K$ UEs. The presented achievable sum rate results are averages of five independent trials.
@@ IV-C Benchmarks for Comparison
Upper Bound: we use the WMMSE-PI algorithm [Jin2024] with 100 iterations, to obtain the optimal AP beamforming matrix $\textbf{W}_{\text{opt}}$ and continuous valued RIS phase shifts $\boldsymbol{\theta}_{\text{opt}}$.
Benchmark 1 (DQNN [Xu2022]): a DL network that quantizes the phase shifts to $\pm\frac{\pi}{2}$, with each quantized value encoded using one bit. Therefore, the resulting control message has length $B=N$. The same DL network is also used to obtain the beamforming matrix based on the CSI [Xu2022, eq. (10), Fig. 2].
Benchmark 2 (ACFNet [Li2024]): a DL network that compresses the bits of the control message. The DNN layer from [Xu2022, Fig. 2] is used as the input of the ACFNet method with $32H$, $16H$, $8H$, $4H$, and $2M+N$ neurons ($H=N+2MK$), in [Li2024, Fig. 2] to obtain the beamforming matrix and phase shift vector before compression. For a fair comparison, compression is set to $B=N_{c}$ control bits by removing the Policy Network in [Li2024, eq. (8) and Fig. 2] and the phase shifts at the RIS are continuous valued by removing $Q_{2}$ in [Li2024, eq. (9) and Fig. 2].
Benchmark 3 (linQ [Shlezinger2019]): a linear quantization method that updates the beamformer and compresses the bits of the control message. Given $\textbf{W}_{\text{opt}}$ and $\boldsymbol{\theta}_{\text{opt}}$, we implement linear analog combining and quantization [Shlezinger2019] (i.e., NNs are matrix multiplication with bias and no activation function) to solve (P1), where $\boldsymbol{\theta}=f_{\text{D}}(f_{\text{Q}}(f_{\text{C}}(\boldsymbol{\theta}_{\text{opt}})))$ and $\textbf{W}=f_{W}(\textbf{W}_{\text{opt}})$. The linear operations ($f_{\text{C}}$, $f_{\text{D}}$, $f_{W}$) compress/decompress the phase shifts and update the beamformer.
Benchmark 4 (AQE): The proposed AQE NN is trained without the WMMSE Beamforming Updater module. The output of this network is: $\textbf{W}=\textbf{W}_{\text{opt}}$ and $\boldsymbol{\theta}=f_{\text{DE}}(f_{\text{Q}}(f_{\text{EN}}(\mathcal{I})))$.
All benchmarks use the same quantization scheme $f_{\text{Q}}(\cdot)$ outlined in Section IV-A to construct the control message. In the DQNN, ACFNet, linQ, and AQE benchmark methods, the beamforming matrix W does not depend on the actual phase shifts $\boldsymbol{\theta}$ used by the RIS controller.
@@ IV-D Results
In Fig. 3, we compare the achievable sum-rate for the proposed method AQE-WMMSE against the benchmarks. In Fig. 3(a), we show the performance when the number of bits is fixed to be the number of RIS elements $B=N_{c}=N=100$. We also present the performance of each method when the control message fails to reach the RIS controller, in which case the RIS uses random phase shifts. Fig. 3(b) and 3(c) compare the compression performance of each method where the number of bits of the control message becomes lower than the number of RIS elements, $B=N_{c}\leq N=100$. We can see that for high transmit power, AQE-WMMSE maintains a high achievable sum rate when compressing down from 100 bits to 40 bits and outperforms other methods even when compressing down to 10 bits.
@@ V Conclusion
A wireless communication informed DL architecture was proposed to jointly compress the RIS phase shifts and update the AP beamforming matrix by reconstructing the phase shift information into a control message (transmitted between the AP and RIS controller). Numerical results showed that updating the beamforming matrix based on the effective channel (which incorporates the actual RIS reflection coefficients) achieves a higher achievable sum rate than recent methods in the literature when the number of bits in the control message is lower than the number of RIS elements.